OpenStudy (anonymous):

Given PO is perpendicular to RS at T, Point O is center Prove: Sum of measure of arcs of PR and QS is 180

OpenStudy (anonymous):



OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Use s = rθ, where s is the length of an arc subtended by rθ, where r is the radius and θ is the central angle. Angle ∠POR and ∠QOS each measure 90 degrees by definition of perpendicular lines. Using the formula s = rθ, Arc PR = (1)*90 Arc QS = (1)*90 The sum of the arcs is 90 + 90 = 180.

ganeshie8 (ganeshie8):

ok there was a typo in ur q

ganeshie8 (ganeshie8):

question should read as :- Given P\(\color{red}{Q}\) is perpendicular to RS at T, Point O is center Prove: Sum of measure of arcs of PR and QS is 180

ganeshie8 (ganeshie8):

check the q and pls let me knw if thats the correct one.. .

OpenStudy (anonymous):

yep @ganeshie8 ur rite

ganeshie8 (ganeshie8):

thank you, since u closed the q, i assume u got the proof :) ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

good :)